OpenStudy (anonymous):

How do you find the tangent line to the graph of "f" at the given point in: f(x)= x^2 + 1

OpenStudy (ddcamp):

Is this a calculus question?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

find derivative.

OpenStudy (anonymous):

it'd be 2x

OpenStudy (anonymous):

I don't understand how you got 2x??

OpenStudy (anonymous):

replace x wiht your equation. later on you'll find a different way of doing so.

OpenStudy (anonymous):

Yes but what about the equation for the tangent line??

OpenStudy (ddcamp):

To find the slope at x, take the derivative: f'(x) = 2x To find the equation of the tangent line, use point-slope form, with (x₁, y₁) = (t, f(t)), and m = 2t: y - (t² + 1) = 2(x - t) y - t² - 1 = 2x - 2t y = 2x + (t² - 2t + 1)

OpenStudy (ddcamp):

Where t is the given x value.